OpenStudy (anonymous):

p= 2L + 2w for L=38 and w=99

OpenStudy (anonymous):

You are given the equation of the perimeter of a rectangle and are asked to find its width given the perimeter and the length. P=2w+2l⟹38=2w+2⋅12⟹w=7.

OpenStudy (anonymous):

P = 2w + 2l P = 38 l = 12 38 = 2w + 2 * 12 38 = 2w + 24 14 = 2w 7 = w

OpenStudy (anonymous):

p=

OpenStudy (anonymous):

they want p=

OpenStudy (anonymous):

P = 38 i think